A DVD has a maximum speed of 7200 revolutions per minute(rpm). If a DVD diameter is 12cm, what is the linear speed in km/h of a point form the disc's center if it is spinning at 5400 rpm?
Answer #1

here,

radius of dvd, r = 12/2 = 6 cm = 0.06 m

final angular speed, w = 5200 * 0.105 = 544.543 rad/s

linear speed = angular speed * radius of path

linear speed = 544.543 * 0.06

linear speed = 32.673 m/s * (3600 s/1000 km)

linear speed = 117.623 km/hr